OpenStudy (anonymous):

Suppose the total cost C(x) in dollars of producing x = 100 units of a product is $2000 and the marginal cost is $2 per unit. At what rate is the average cost C(x) of the items changing with respect to x (the marginal average cost)?

OpenStudy (anonymous):

in this question do you mean C(100) = 2000$ or C(100) = 2000 +2 x 100 ?

OpenStudy (anonymous):

i'm pretty sure c(100) = 2000$ is what is meant....it is on my review sheet so i'm not entirely sure....

OpenStudy (anonymous):

what does marginal cost mean here?

OpenStudy (anonymous):

hang on...i may have figured it out, give me like 5 min

OpenStudy (anonymous):

okay i don't think i did...marginal cost is the rate of change with respect to x: C'(x)=marginal cost function the average cost function is Cbar(x) = c(x)/x and Cbar'(x) = marginal avg. cost function

OpenStudy (anonymous):

okay maybe marginal cost is the chng in rate of of product and if these goes like this then according to your question C(x) = 2000 right so avg cost for whole product is 2000/100= 20 total marginal cost fr 100 products = 200 so actual cost is (2000 - 200)/100= 1800/100 = 18 per product

OpenStudy (anonymous):

i got this much but i'm not able to understand last lines in question ....

OpenStudy (anonymous):

it's asking for the marginal average cost and the answer is Cbar'(100)= -9/50....but i don't know how to get that

OpenStudy (anonymous):

u must be frm economics ? hun .... never came acrs this type of problems .. explain me little abt each term.. and i'l get u ur answer..

OpenStudy (anonymous):

but u must see that if we divide 18/100(avg actual cost )/(total products) we are able to get wt ur question demand

OpenStudy (anonymous):

wait c(x) = 18 + 2x

OpenStudy (anonymous):

it's calculus but very related to economics obviously...i don't know enough about it to help you anymore....this question has ver yconfused! i really appreciate your help, hopefully it just a poorly worded question (or something)...i will skip this one and ask about it first thing in the morning

OpenStudy (anonymous):

ohhh....how did you find that?!

OpenStudy (anonymous):

so c(x)/x = 18/x + 2

OpenStudy (anonymous):

well 18 is the real cost .... and 2 marginal so it incrs with every product so c(x) = 18 +2x

OpenStudy (anonymous):

ohhhh yes!!

OpenStudy (anonymous):

now u said about c'(X) ! prvsly

OpenStudy (anonymous):

so Cbar'(x) which is the marginal avg. cost is -18x^-2

OpenStudy (anonymous):

wait make c(x) = 1800 + 2x now u'l get ur answer !

OpenStudy (anonymous):

c'(x)=-1800/x^2=-1800/10000=-18/100=-9/50

OpenStudy (anonymous):

did u undrstnd it now?

OpenStudy (anonymous):

how did you make 1800+2x get to -1800/x^2?

OpenStudy (anonymous):

c(x)= 1800 + 2x c(x)/x = 1800/x +2

OpenStudy (anonymous):

wt abt now?

OpenStudy (anonymous):

yep!!! i got it! thank you sooooo much!!!!!!!! you worked hard!

OpenStudy (anonymous):

you too !!